Please show calculations. Thank you!

PinkCo, Inc., operates in two states, both of which equally weight the three apportionment factors. PinkCo reports the following results for the year. Amounts are stated in millions of dollars.

State A State B Totals
Sales $25 $75 $100
Payroll 20 30 50
Property 0 100 100

Compute the apportionment percentage for both states.

Round your answer to two decimal places. Then, enter it as a whole number and not a decimal. For example, 0.4566 would be rounded to 0.46 and entered as 46.

The apportionment percentage for State A is _______% and for State B is _______%.
Answer #1

Calculate apportionment percentage

State A State B Total
Sales 25 75 100
Payroll 20 30 50
Property 0 100 100
Total 45 205 250

Apportionment percentage of State A = 45/250 = 0.18

State B = 205/250 = 82%

The apportionment percentage for State A is 18% and for State B is 82%.